Network Solutions just introduced a​ new, fully automated manufacturing plant that produces 1,500

wireless routers per day with materials costs of $50 per router and no other costs. The average number of days a router is held in inventory before being sold is 54 days. In​ addition, they generally pay their suppliers in 27 ​days, while collecting from their customers after 26 days.

a.  What is the cash conversion​ cycle?

b.  What would happen to the cash conversion cycle if they could stretch their payments to suppliers from 27

days to 47 ​days?

c.  How much would working capital be reduced if they stretched their payments to suppliers from 27 days to 47

​days?
Answer #1

Please find answers to first two sub parts below

a)Cash Conversion Cycle = Days Inventory Oustanding(DIO) + Days Sales Outstanding(DSO) - Days Payable Outstanding(DPO) ------ 1

Total Number of wireless routers produced in a day = 1500

Cost of Materials per router = $50

Average Number of Days Router held in Inventory before being Sold( DIO - Days Inventory Outstanding) = 54 days ------2

Average number of days to pay back suppliers (DPO - Days Payable Outstanding) = 27 days ----- 3

Average number of days to collect money from customers (DSO - Days Sales Outstanding) = 26 days ------ 4

(Substituting 2,3 and 4 in 1)

Cash Conversion Cycle = 54 + 36 - 27 = 63 days

b) All else remaining constant if payments to suppliers (DPO) is stretched to 47days

Cash Conversion Cycle = 54+36 - 47 =  43 days (Cash Conversion cycle reduces by 20 days)